# Mechanistic Unlearning: Robust Knowledge Unlearning and Editing via Mechanistic Localization

### v1

@@ 2.3 Parameter Update Methods
where $L_{\text{forget}}$ is an unlearning loss on the $D_{forget}$ subset of facts we want to forget, $L_{\text{retain}}$ is a cross-entropy loss on the remaining facts, and $L_{\text{SFT}}$ is a cross-entropy loss on the Pile dataset (Gao et al., 2020). The unlearning loss we use is the $\log(1-p)$ measure from Mazeika et al. (2024), for its stability and fewer side effects.
For editing, we use a loss function $L=\lambda_{1}L_{\text{injection}}+\lambda_{2}L_{\text{retain}}+\lambda_{3}L_{\text{SFT}}$, where $L_{\text{injection}}$ is a cross-entropy loss on the forget facts maximizing the probability of the alternative false target. Our $\lambda$s are in Section A.4.
Weight Masking
For weight masking-based unlearning, we train a binary differentiable mask over individual weights of the model within the localized components, inspired by weight pruning/masking work (Bayazit et al., 2023; Panigrahi et al., 2023). In this case, no weight updates are being performed. Rather, the mask turns a subset of the weights to zero. To control the sparsity penalty, we additionally include $\lambda_{4}*L_{\text{reg}}$, an L1 regularization term.
3 Unlearning and Editing Results
In this section, we show the results of unlearning and editing across all of the mentioned localization techniques for localized fine-tuning and weight masking. We try three unlearning/editing goals: unlearning all athletes playing basketball (referred to as unlearning sports), editing the model’s associations for a constant set of 16 athletes across all sports (akin to the error injection setup from (Hase et al., 2023), and editing the associations for 16 facts from the CounterFact dataset Meng et al. (2023). All the techniques are assessed based on standard and adversarial evaluations.
3.1 Localized Finetuning
3.1.1 Standard Evaluation
For the sports dataset, following Nanda et al. (2023), we first evaluate the accuracy of our models to complete the prompt, “Fact: [athlete] plays the sport of”, with a one-shot example of Tiger Woods playing golf given first. Note that this is the same prompt used to train the unlearning in the first place. We refer to this accuracy as Normal Accuracy. Inspired by Patil et al. (2023) and Lynch et al. (2024), we also use an alternative input and output prompting setup to measure if our unlearning has “overfitted” to the prompt input and the output format. We instead use a multiple-choice format with the choices of football, baseball, basketball, and golf. We refer to the accuracy of the model answering with the ground truth on this prompt format as the MCQ Accuracy (for both unlearning and editing, stronger methods should decrease the MCQ accuracy).
For CounterFact, we evaluate the accuracy of our models to complete the original factual associations. We report the original robustness and side effect evaluations from the Meng et al. (2023) dataset, the Paraphrase and Neighborhood facts accuracies. Since the Paraphrased prompts have the same answer string as the original question, we hypothesize that they may be easier for model editing to generalize to - we want to test that the knowledge is edited under any expression of the knowledge. We perform a knowledge robustness check by phrasing the associations in multiple choice format (MCQ Accuracy), with the true answer, the injected false answer, and two other question-specific LLM-generated incorrect answers. Once again, accuracy is measured with the ground truth answer, so we always want the MCQ accuracy to decrease throughout editing.
Finally, we also evaluate our models’ accuracy on MMLU (Hendrycks et al., 2021) as a proxy for the general side effects of unlearning unrelated to sports and CounterFact. Our results with localized fine-tuning are shown in Table 1 for sports, Table 2 for athletes, and Table 3 for CounterFact.
@@ 3.1.1 Standard Evaluation
Table 2: Localized fine-tuning accuracy on standard evaluations: editing a constant 16 athlete subset, considering accuracy relative to the original ground truth answers
Attrib. Patching 0.447 0.998 0.895 0.612
Causal Tracing 0.586 0.994 0.945 0.613
Manual 0.001 0.970 0.108 0.611
Random 0.883 0.988 0.875 0.614
All-MLPs 0.001 0.965 0.166 0.574
Nonlocalized 0.354 0.890 0.155 0.573
Table 3: Localized fine-tuning accuracy from editing 16 facts from the CounterFact dataset, considering accuracy relative to the original answers
Localization Type Forget $\downarrow$ Maintain $\uparrow$ MCQ $\downarrow$ MMLU $\uparrow$ Paraphrase $\downarrow$ Neighborhood $\uparrow$
Attrib. Patching 0.000 0.974 0.944 0.691 0.286 0.681
Causal Tracing 0.000 0.973 0.481 0.692 0.032 0.736
Manual 0.000 0.965 0.295 0.691 0.035 0.733
Random 0.000 0.979 0.651 0.690 0.159 0.757
All-MLPs 0.000 0.927 0.390 0.688 0.083 0.725
Nonlocalized 0.000 0.920 0.411 0.682 0.104 0.746
As seen from the tables, across all tasks, unlearning with manual localization achieves the highest robust multiple-choice forget accuracy and near the highest MMLU, and very competitive normal forget and retain accuracy. Only manual localization, all-MLPs, and nonlocalized approaches had generalized their unlearning to the multiple choice format, but manual localization has higher MMLU performance and higher retain accuracy than All-MLPs and nonlocalized. On the CounterFact task, only manual interpretability achieves baseline MCQ accuracy, and manual localization also achieves the second-lowest score on the Paraphrase robustness evaluation as well. As before, because the Paraphrased facts may share the same attribute extraction mechanisms to the output as the original facts, we emphasize that the MCQ evaluation is the higher signal robustness evaluation.
This indicates that OT localization methods do not robustly modify the model, and the supposedly-unlearned/edited information can be extracted through prompt and task variations.
3.1.2 Adversarial Relearning
We measure the ability of our models to withstand adversarial relearning, both to address the scenario in which adversaries may have fine-tuning access and as an upper-bound measure for the quality of unlearning – a model that needs relatively fewer steps to relearn may not have properly unlearned [^1] Ideally, one would compare here to the number of iterations needed to learn these facts for a “gold-standard” model that was never trained on these facts during pre-training. Obtaining the gold-standard model is prohibitively expensive. Therefore, we rely on a heuristic that unlearned models that need longer re-learning behave more like the gold-standard one.. We retrain with a rank-64 LoRA across all linear modules using a subset of the forget dataset for sports, with details available in Section A.5.2. We measure the ability of the relearned model to answer with the ground truth statements, similar to the methodology of Deeb & Roger (2024).
Figure 2 (left) and Figure 2 (right) compare relearning robustness of different localization techniques for sports and athletes respectively.
As shown in Figure 2 (left), for sports-unlearned models, manual interpretability is the localization method that is most robust to the low-resource relearning. Unlearning based on every other localization as well as the no-localization technique regains accuracy on the rest of the forget set within a few iterations. For all of the athlete-edited models, relearning on just two of the edited athletes recovers significant accuracy on all athletes in all except for the manual interpretability model.
Figure 2: (Left) Retraining basketball-unlearned models with two athletes in the forget set, for ten iterations. (Right) Retraining models that have had 16 athlete-sport associations replaced with Golf as the athletes’ associated sport, with two athletes in the forget set, for ten iterations. The y-axis represents the normal accuracy on the forget set. Low-resource relearning of athletes demonstrates that using manual interpretability makes edits significantly more robust to relearning.
For CounterFact relearning, we change the setup slightly: we retrain on all forgotten statements, since the set of edited statements aren’t as directly related as for sports facts. We measure accuracy using the same train set of edited/forgotten statements in Figure 3 (left). However, because we are directly optimizing the accuracy on this train set, we also try the multiple choice evaluation on this train set to see if the relearning generalizes (if the model has reinternalized this knowledge), in Figure 3 (right). Figure 3 (right) demonstrates that the correct factual associations were not deeply relearned in any of the manual interpretability, nonlocalized, or all MLP models, with the manual interpretability model having the lowest MCQ accuracy.
Figure 3: Retraining counterfact-edited models with all forgotten facts, for twenty iterations. (Left) The y-axis represents the normal accuracy on the forget set. (Right) The y-axis represents the Multiple Choice Formatted question accuracy on the forget set.
3.1.3 Latent Knowledge Analysis
In this section, we provide more evidence of our hypothesis that FLU unlearning targets the source of intermediate latent knowledge rather than the extraction mechanism. We use the Sports Facts dataset for this study, as the answer is always one of three representations (the various sports).
Similar to Patil et al. (2023), we train logistic regression models (probes) (Alain & Bengio, 2018) on the activations of every model layer to predict the correct ground truth sport from the prompt, with the idea that a model that has truly modified a fact would not have much predictive value in its activations. For more details on probe training, see Section A.5.3.
We test whether the models retain information about the original athlete set in the intermediate layer representations after editing. Figure 4 shows the accuracy of the trained per-layer probes on the set of edited athletes. If model updates were concentrated in the extraction mechanism, we should still see high probe accuracies, as the probes are accessing the latent representations.
Figure 4: Probe accuracy (combined over all three sports) on the athlete golf-injected models across layers. The ‘Localized AP’, ‘Localized CT’, and ‘Random’ lines are overlapping.
Figure 4 provides evidence that the manual interpretability, All-MLPs, and nonlocalized athlete-edited models contain less recoverable representations of the edited associations in the intermediate layers, with the probes having the least accuracy on the manual interpretability model. This suggests our editing guided by manual interpretability has properly interfered with the lookup MLPs enriching the latent stream with information about the correct answer.
3.2 Weight Masking
@@
In this section we employ weight masking to quantify the size of weight updates needed to unlearn/edit facts, for more direct comparisons. We empirically evaluate how a learned binary mask over individual weights of the localized components can produce editing/unlearning, and vary the size of this mask.
After weight masking, we also analyze the proportion of each mechanism (fact lookup, attribution extraction) that is masked for each localization in Section A.2.3. We demonstrate that OT methods and nonlocalized editing all modify a higher proportion of the extraction head/MLP parameters than the fact lookup mechanism parameters, supporting our claim that OT methods target extraction mechanisms rather than the fact lookup mechanisms needed for robustness.
3.2.1 Standard Evaluation
@@
For editing the subset of athletes, Figure 7 shows that causal tracing localization causes the model to have 0% accuracy on the forget set, and manual interpretability and nonlocalized editing cause the model to have near guessing rate (33%) accuracy. However, only manual localization minimizes loss of capabilities while editing the athlete subset (Figure 8).
@@
We find similar results for editing on the CounterFact dataset. However, we find minimal difference in MMLU accuracy in all methods at all numbers of masked weights. Thus, we instead report the maintain and forget accuracies of these facts at different discretization thresholds in Figure 9. We provide additional performance metrics in Section A.3.
@@
4 Discussion
Our work evaluates the efficacy of different localization methods for modifying factual associations. We demonstrate clear benefits of localization for unlearning robustness through localized fine-tuning combined with manual mechanistic interpretability techniques designed for fact recall.
In Section 3.1.3 and Section A.2.3, we provide evidence that OT and baseline approaches fail to be robust because they target easily-localizable and high direct logit importance attention head components, that transform existing latent factual knowledge to the desired output format. This can fail to generalize to different input and output formats and does not target the true source of knowledge in the model: other input/output formats can allow alternative attention mechanisms to transform this knowledge, and low-resource relearning can quickly repair the original attention mechanism. In contrast, FLU mechanistic understanding allows us to target unlearning at the sites where knowledge is sourced, which robustly prevents that information from entering the latent stream in any format.
Acknowledgements
We thank Stephen Casper, Aengus Lynch, and Max Li for their advice with unlearning methods and interpreting results. We also thank Daniel M. Roy, Amr Khalifa, and Eleni Triantafillou for feedback on various drafts of this work. This project used compute provided by the Center for AI Safety.
Appendix A Appendix
A.1 Fact Injection Results
We additionally consider the common factual editing methodology, specifically an error injection setup (Hase et al., 2023) where we replace correct athlete-sport associations with incorrect associations between athlete and ”Golf”. The results are in Table 4 and Table 2, where we demonstrate that our manual localization method again achieves the strongest robust unlearning generalization while maintaining more general capabilities than the other robustly unlearned model.
Table 4: Results of unlearning basketball associations, with the objective of replacing the correct sport of ”Basketball” with ”Golf”. Forget refers to the model’s accuracy at stating the original sport association which should have been replaced.
Attrib. Patching 0.000 1.000 0.815 0.611
Causal Tracing 0.028 1.000 0.866 0.614
Manual 0.035 0.973 0.257 0.610
Random 0.018 1.000 0.839 0.611
All-MLPs 0.000 0.946 0.363 0.571
Nonlocalized 0.000 0.995 0.376 0.565
Table 5: Localized fine-tuning accuracy on standard evaluations: Unlearning a constant 16 athlete subset, retaining all other facts.
Attrib. Patching 0.941 0.964 0.934 0.614
Causal Tracing 0.891 0.915 0.910 0.612
Manual 0.034 0.975 0.175 0.615
Random 0.938 0.952 0.883 0.612
All-MLPs 0.003 0.973 0.281 0.599
Nonlocalized 0.203 0.570 0.391 0.540
We also perform relearning and latent knowledge experiments in Figure 2 (right) and Figure 4 (Appendix A), demonstrating that manual localization for the athlete subset injection improves relearning and latent knowledge robustness.
Figure 10: Retraining athlete-unlearned models with two athletes in the forget set, for ten iterations. The y-axis represents the normal accuracy on the forget set. Low-resource relearning of athletes demonstrates that manual and non-localized unlearning techniques are robust to this test (staying close to the guessing rate of 33%), while the other methods maintain full performance on the entire forget set.
A.2 Gemma Interpretability Analysis
A.2.1 Sports Facts
We redo analysis from Nanda et al. (2023) on Gemma-7B. We train logistic regression models (”probes”) to predict the correct sport given the internal representation of the model at a layer. We find that probes predicting the correct sport increase in accuracy significantly in layers 2 through 7, and we find the mean ablation of all attention heads past layer 7 to have minimal impact on the linear representation of player attributes (Figure 12).
Unlike Nanda et al. (2023), however, we find attention heads past layer 2 that impact the linear representation of attributes and thus could potentially be important for fact lookup (Figure 12). However, because they could likely play a variety of other different roles such as token concatenation, following the findings of Geva et al. (2023); Nanda et al. (2023) that MLPs do primary factual representation enrichment, in this work we only consider the MLPs as our localization.
Figure 11: Difference in final layer probe accuracy when mean ablating a single head for all heads between layers 0 and 6.
Figure 12: Probe accuracy on predicting sport across layers. ”Prompt” refers to the entire facts prompt, while ”Athlete” is just the athlete’s name.
A.2.2 CounterFact
We repeat analysis from Nanda et al. (2023) and Geva et al. (2023) on Gemma-2-9B. We first measure the effect on the difference in logits between correct and incorrect answers of facts when patching the direct path of attention heads and MLPs to the final output, shown in Figure 14 and Figure 14. An attention head or MLP will have a large effect on the logit difference if it is important in moving the factual information to the last token position or decoding it into the correct answer. We call these components part of the ”fact extraction mechanism”, and aim to find the source of the factual information moved by this mechanism.
To find this source, we patch the outputs of MLPs to this ”fact extraction mechanism” and measure the resultant change in logit difference (Figure 15). An MLP would cause a large change in logit difference if it caused relevant representations to form that are then moved by the ”fact extraction heads” to increase the probability of the correct output. We take the MLPs with the highest change and include them in our manual localization of CounterFact (MLPs 3-5, 7-10, 14-17).
Figure 13: Change in logit difference when patching heads to the final output.
Figure 14: Change in logit difference when patching MLPs to the final output.
Figure 15: Change in logit difference when patching MLPs to the extraction mechanism.
A.2.3 Mechanism Weight Analysis

### v2

@@ A.1 Sports Unlearning Results
where $L_{\text{forget}}$ is an unlearning loss on the $D_{forget}$ subset of facts we want to forget, $L_{\text{retain}}$ is a cross-entropy loss on the remaining facts, and $L_{\text{SFT}}$ is a cross-entropy loss on the Pile dataset (Gao et al., 2020). The unlearning loss $L_{\text{forget}}$ we use is the $\log(1-p)$ measure (where p is the probability of the correct sport) from Mazeika et al. (2024) due to its empirical stability and fewer side effects: vanilla gradient ascent more strongly incentivizes the model to have significantly lower logprobs than wouldn’t be encountered in a model that has not been trained on the factual association, and it detracts from model maintenance of $L_{\text{retain}}$ and $L_{\text{SFT}}$.
We present results on using various localizations on the Sports-Unlearning task in Table 1. The FLU localization allows unlearning to be more robust to the MCQ prompt format while maintaining performance on the MMLU dataset.
@@
A.2 FLU Interpretability Analysis
A.2.1 Sports Facts
We redo analysis from Nanda et al. (2023) on Gemma-7B, Gemma-2-9B, and Llama-3-8B. We train logistic regression models (”probes”) to predict the correct sport given the internal representation of the model at a layer. We find that probes predicting the correct sport increase in accuracy significantly in layers 2 through 7 in Gemma-7B and 2 through 8 for Gemma-2-9B and Llama-3-8B (Figure 8).
Unlike Nanda et al. (2023), however, we find attention heads past layer 2 that impact the linear representation of attributes and thus could potentially be important for fact lookup. However, because they could likely play a variety of other different roles such as token concatenation, following the findings of Geva et al. (2023); Nanda et al. (2023) that MLPs do primary factual representation enrichment, in this work we only consider the MLPs as our localization.
Figure 8: Accuracy of probes predicting the correct sport across layers for different models.
A.2.2 CounterFact
We repeat analysis from Nanda et al. (2023) and Geva et al. (2023) on Gemma-2-9B. We first measure the effect on the difference in logits between correct and incorrect answers of facts when patching the direct path of attention heads and MLPs to the final output, shown in Figure 9. An attention head or MLP will have a large effect on the logit difference if it is important in moving the factual information to the last token position or decoding it into the correct answer. We call these components part of the ”fact extraction mechanism”, and aim to find the source of the factual information moved by this mechanism.
To find this source, we patch the outputs of MLPs to this ”fact extraction mechanism” and measure the resultant change in logit difference (Figure 10). An MLP would cause a large change in logit difference if it caused relevant representations to form that are then moved by the ”fact extraction heads” to increase the probability of the correct output. We provide the logit differences for all 64 facts along with just the first 16 facts, and see that the logit differences are similar across the dataset splits. We take the MLPs with the highest change ($>0.02$) and include them in our FLU localization of CounterFact.
Figure 9: Change in logit difference when patching attention heads to the output.
Figure 10: Logit difference when patching MLPs to the extraction mechanism found above for different models.
A.3 OT Selected Components
What MLPs do the automated OT methods localize? We explore the attribution scores of the automated localization methods (causal tracing and attribution patching) on the MLPs to see if automated localization methods can detect the FLU mechanism. In Figures 11 and 14, for Gemma-7B, we see that both CT and AP localizations target the later layer MLPs instead of the FLU mechanism (Figure 11).
Figure 11: Attribution scores on MLPs on sports facts for Gemma-7B.
For AP localization, this trend continues with Gemma-2-9B (Figure 12) and Llama-3-8B (Figure 13). However, CT localization does highlight some of the early layer MLPs that are in the FLU mechanism, especially for Gemma-2-9B.
Figure 12: Attribution scores on MLPs on sports facts for Gemma-2-9B.
Figure 13: Attribution scores on MLPs on sports facts for Llama-3-8B.
We repeat this analysis on CounterFact in Figures 14, 15 and 16. Again we see AP localizations in particular assign higher scores to later-layer MLPs, and CT only highlights FLU components on Gemma-2-9B, localizing other extraction layers on the other models.
Figure 14: Attribution scores on MLPs on CounterFact facts for Gemma-7B.
Figure 15: Attribution scores on MLPs on CounterFact facts for Gemma-2-9B.
Figure 16: Attribution scores on MLPs on CounterFact facts for Llama-3-8B.
A.4 Weight Masking
@@
In this section we employ weight masking to quantify the size of weight updates needed to unlearn/edit facts, for more direct comparisons. Our loss function $L=\lambda_{1}L_{\text{forget}}+\lambda_{2}L_{\text{retain}}+\lambda_{3}L_{\text{SFT}}$ + $\lambda_{4}L_{\text{reg}}$ now includes an L1 regularization term to control the sparsity. We empirically evaluate how a learned binary mask over individual weights of the localized components can produce editing/unlearning, and vary the size of this mask. “Manual Interp” refers to the FLU localization technique for all the following results in this section.
A.4.1 Unlearning Sports
@@
A.4.2 Editing Athletes
For editing the subset of athletes, Figure 19 shows that causal tracing localization causes the model to have 0% accuracy on the forget set, and FLU and nonlocalized editing cause the model to have near guessing rate (33%) accuracy. However, only manual localization minimizes loss of capabilities while editing the athlete subset (Figure 20).
@@ A.4.2 Editing Athletes
A.4.3 Editing CounterFact
We find similar results for editing on the CounterFact dataset. However, we find minimal difference in MMLU accuracy in all methods at all numbers of masked weights. Thus, we instead report the maintain and forget accuracies of these facts at different discretization thresholds in Figure 21.
@@ A.4.3 Editing CounterFact
Additionally, we report a comparison of all localizations across discretization thresholds for normal and MCQ forget sets in Figure 23 and Figure 23. We see that FLU outperforms all other methods of localization in preserving maintain accuracy while decreasing forget accuracy.
We perform additional adversarial analysis of accuracies across different discretization thresholds. We report the ”paraphrase” and ”neighborhood” adversarial results in Figure 25 and Figure 25, but find no significant results.
A.5 Mechanism Weight Analysis